# High-flux cold lithium-6 and rubidium-87 atoms from compact two-dimensional magneto-optical traps

### v1

@@ I Introduction
Ultracold polar molecules have emerged as a powerful platform for studying ultracold chemistry, testing fundamental symmetry, and exploring strongly correlated many-body systems with strong dipole-dipole interactions [1, 2]. Among bialkali polar molecules, 6Li87Rb ground-state molecules stand out as they exhibit the second-largest electric dipole moment [3]. This makes it a promising candidate for realizing and probing novel quantum phases, such as the long-sought dipolar Bardeen-Cooper-Schrieffer to Bose-Einstein condensate (BCS-BEC) crossover [4]. So far, several labs have successfully produced degenerate mixtures of fermionic lithium and bosonic rubidium atoms [5, 6] and observed Feshbach resonances between lithium and rubidium atoms [5, 7]. Ultracold 6Li87Rb polar molecules remain elusive in experiments.
@@
Next-generation experiments with ultracold dual-species atoms and polar molecules require shorter cycle times and miniaturized system volume, thereby demanding a compact dual-species cold-atom source with high atomic flux. In the early days, achieving high flux often necessitated complex setups with multiple bulky vacuum chambers, especially for high-melting-point elements such as lithium and sodium [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Those elements typically require long Zeeman slowers to decelerate fast atomic beams from hot ovens. While compact, two-dimensional magneto-optical traps (2D MOTs) have relatively low capture velocities, therefore tend to work more efficiently for low-melting-point elements such as rubidium and cesium [23, 24, 25, 26, 27, 28]. In 2009, Tiecke et al. demonstrated a compact lithium 2D MOT using permanent magnets, achieving loading rates comparable to those of Zeeman slowers [29]. Recently, a modularized lithium quantum gas setup based on a 2D MOT has been developed in Heidelberg and Munich [30, 31], which further enhances the loading rate by employing high-power and top-hat cooling beams [30]. Combining a 2D MOT with a short-distance Zeeman slower can substantially enhance its capture velocity and thereby loading rates. This has been demonstrated for sodium, strontium, and ytterbium atoms [32, 33, 34, 35, 36], but not for lithium atoms yet. Moreover, integrating high- and low-melting-point elements into a single compact system while maintaining high flux for both species remains an experimental challenge.
@@
In this work, we present a compact and high-flux dual-species cold-atom source that overcomes these challenges by combining a rubidium 2D MOT in series with a lithium 2D MOT, enhanced by a short-distance Zeeman slower for lithium. As a result, the 3D MOT loading rate of lithium atoms reaches a record value that is 44 times higher than that without the Zeeman slowing light. The flux of rubidium also reaches a reasonably high value. Meanwhile, the vacuum chamber is so small that it can be mounted on a movable breadboard [30]. Thereby, we can smoothly extract it from the magnetic coils and the forest of optics for laser beam alignment or maintenance, then push it back to the original position. Thanks to the in-series 2D MOTs configuration, the compact vacuum system also allows more than 270 degrees wide optical access to the science cell in the horizontal plane, providing an excellent starting point for future experiments.
@@ II.1 Vacuum system and general setup
Our vacuum system integrates Rb and Li two-dimensional magneto-optical traps (2D MOTs) arranged in series inside a high-vacuum (HV) region. These sources generate high-flux pre-cooled atomic beams that are further collimated by push beams and captured in a dual-species 3D MOT located in an ultra-high-vacuum (UHV) science cell, as shown in Fig. 1. Two differential pumping channels (DP) maintain a pressure drop of more than three orders of magnitude between the HV and UHV regions while preserving large solid angles for both beams (see Appendix A). Placing both 2D MOTs on the same side of the main chamber maximizes optical and microwave access to the science cell for manipulation of ultracold atoms and molecules [37]. During daily operation, we measure a pressure on the order of $1\times 10^{-9}\ \mathrm{mbar}$ in the Li octagon chamber and $3\times 10^{-11}\ \mathrm{mbar}$ in the cubic science chamber. This pressure is consistent with the measured vacuum-limited lifetime exceeding 50s50\ $\mathrm${s} in the science chamber.
The Rb 2D MOT loads atoms directly from the background vapor at a pressure of about $10^{-8}\ \mathrm{mbar}$ and provides an atomic flux of roughly $10^{10}\ \mathrm{atoms/s}$ in a $2\mathrm{D}^{+}$ configuration. In contrast, $^{6}\mathrm{Li}$ has an extremely low vapor pressure at room temperature ($\sim 10^{-20}\ \mathrm{mbar}$[38] and therefore requires an oven-based source. We place the high-vapor-pressure, low-divergence Rb 2D MOT at the upper stream of the vacuum system, while the low-vapor-pressure, high-divergence Li 2D MOT is located in the middle, with the crucial UHV science chamber further downstream. The Li 2D MOT chamber also serves as an additional differential pumping stage between the Rb 2D MOT region and the science chamber. This layout facilitates a natural HV-to-UHV transition, allowing for the independent optimization of the spatially separated Rb and Li 2D MOTs, thereby enabling the simultaneous high-flux loading of both species into the 3D MOT.
@@ II.2 Source and Ovens
The Rb source (CAS: 7440-17-7, CAT: 044214.03, Thermo Fisher Scientific Inc.) consists of 1 gram of natural isotopic abundance rubidium ($\eta_{{}^{87}\mathrm{Rb}}=27.83(2)\%$ and $\eta_{{}^{85}\mathrm{Rb}}=72.17(2)\%$ [40, 41]), sealed in a glass ampoule to prevent oxidation. We designed the Rb oven using a copper tube with a specialized opening that serves as a gasket (LIMIT VACUUM, Beijing). This assembly is sealed with a custom flange and connected to the Rb chamber via an angle valve, allowing for a controlled and gentle introduction of the atomic vapor. To activate the Rb source, we mechanically crush the copper tube, breaking the enclosed ampoule during the final stage of the vacuum baking process.
@@ II.3.1 Lithium laser system
The $^{6}\text{Li}-D_{2}$ transition possesses a natural linewidth of $\Gamma_{\mathrm{Li}}/2\pi\approx 5.87\ \mathrm{MHz}$, while the hyperfine splitting between the upper manifold $|2^{2}\text{P}_{3/2};F^{\prime}=1/2,3/2,5/2\rangle$ is merely 4.4MHz4.4\ $\mathrm${MHz} [42, 38]. While driving the cycling transition $|2^{2}\text{S}_{1/2};F=3/2\rangle\to|2^{2}\text{P}_{3/2};F^{\prime}=5/2\rangle$ for laser cooling, the transitions to the $|2^{2}\text{P}_{3/2};F^{\prime}=1/2,3/2\rangle$ states are simultaneously driven. Consequently, no convenient fully closed cooling transitions can be utilized here [29].
@@
Figure 2: Laser frequency scheme (not to scale). Left: $^{6}\mathrm{Li}$ $D_{2}$ transition from the $2^{2}\mathrm{S}_{1/2}$ state to the $2^{2}\mathrm{P}_{3/2}$ manifold. Red arrows indicate the corresponding laser frequencies, and $\delta^{\mathrm{Li}}_{i}$ denote the detunings of the individual beams with respect to the upper manifold. Transition frequencies are taken from [38]. Right: $^{87}\mathrm{Rb}$ $D_{2}$ transition from the $5^{2}\mathrm{S}_{1/2}$ state to the $5^{2}\mathrm{P}_{3/2}$ state. Red arrows indicate the corresponding laser frequency, and $\delta^{\mathrm{Rb}}_{i}$ denote the detunings of the individual beams with respect to the corresponding transition. Transition frequencies are taken from [40].
@@ II.3.2 Rubidium Laser System
The laser frequency scheme for laser cooling of $^{87}\text{Rb}$ is shown in Fig. 2, following many of the well-established Rb cooling schemes, for example [43]. In contrast to $^{6}\text{Li}$, the $^{87}\text{Rb}\ D_{2}$ transition exhibits a natural linewidth of $\Gamma_{\mathrm{Rb}}/2\pi\approx 6.07\ \mathrm{MHz}$, with hyperfine splittings in the $5^{2}\mathrm{P}_{3/2}$ states ranging from tens to hundreds of $\mathrm{MHz}$ [40, 44]. These well-resolved energy separations significantly simplify the implementation of laser cooling. For our cooling scheme, we utilize the closed $|5^{2}\mathrm{S}_{1/2};F=2\rangle\to|5^{2}\mathrm{P}_{3/2};F^{\prime}=3\rangle$ transition as the primary cooling cycle, supplemented by a repumping beam resonant with the $|5^{2}\mathrm{S}_{1/2};F=1\rangle\to|5^{2}\mathrm{P}_{3/2};F^{\prime}=2\rangle$ transition to recover atoms that decay out of the cooling cycle. Analogous to our lithium configuration, we implement a push beam, red-detuned from the $|5^{2}\mathrm{S}_{1/2};F=2\rangle\to|5^{2}\mathrm{P}_{3/2};F^{\prime}=3\rangle$ transition, to enhance the directed atomic flux along the axial direction.
@@ II.5 Li short-distance Zeeman slowing design
The simulations were conducted based on the simplified models outlined in [29, 32, 35]. The total atomic flux emitted through the oven aperture (area $A$) is given by:
@@
This formula is a good approximation for the capture velocity $v_{c}\ll\alpha$[29]. Here, $a_{6}$ is the isotopic abundance of $^{6}\text{Li}$, $\alpha=\sqrt{2k_{B}T/m}$ is the most probable velocity, and $\Omega_{c}$ is the solid angle subtended by the 2D MOT region. The oven temperature $T$ determines the velocity distribution and the saturated vapor pressure $p_{s}$, which is derived from the Clausius-Clapeyron relation. Further details can be found in [29].
Figure 4: Simulation of the short-distance Li Zeeman slowing and 2D MOT. (a) Velocity distribution of $10^{6}$ Li atoms with and without Zeeman slowing. The Zeeman slower reshapes the initial distribution by decelerating atoms in the blue region into the orange peak, bringing them below the 2D MOT capture velocity $v_{c}$. Inset: velocity deformation caused by Zeeman slowing (figure inspired by [32]). (b) Numerical simulations of atomic trajectories along the applied magnetic field. The 2D MOT beam region is shown in orange. The light gray area indicates the capture velocity $v_{c}$ of the 2D MOT.
@@
To minimize contamination of the glass viewport caused by lithium’s high sticking coefficient and small mass, the Zeeman beam is redirected using a gold-coated in-vacuum mirror mounted 300mm300\ $\mathrm${mm} above the center of the octagon chamber on the top of a so-called double-T structure (compare Fig. 1). However, we still observe some contamination after several months of operation, although the ZS performance is barely affected. An alternative approach would be to heat a viewport made of sapphire crystal positioned directly against the atomic flux, as demonstrated by Nosske et al. [35]. Their implementation also showed degradation in optical transmission over a period of several months. The ZS beam is far-detuned (approximately $85\Gamma$) to take advantage of the positive gradient region of the magnetic field ($\partial B_{y}/\partial z>0$), minimizing interference with the 2D MOT operation. In contrast to conventional ZS designs [8], our system employs a quadrupole magnetic field oriented transverse to the atomic beam trajectory. This configuration demands circularly polarized light relative to the local magnetic field vector. Consequently, when using a linearly polarized ZS beam (as in our setup), only the $\sigma^{-}$ polarization component effectively contributes to the slowing process, reducing the practical intensity of the ZS beam by 50% [35]. Along one dimension, the radiation force acting on the atoms is described by:
@@
Here, $\delta_{0}$ is the zero-field detuning, $v$ is the atomic velocity, $B$ is the magnetic field strength, and the factor $0<\eta_{B}<1$ accounts for shielding effects of the metal chamber to the magnetic field[32].
@@ II.6 Push Beam
The atom cloud generated by the 2D MOT forms a needle-shaped distribution aligned along the longitudinal axis. This two-dimensional cooling process produces slow atomic beams with velocities typically below 30m/s30\ $\mathrm${m/s} [23]. The push beam serves to accelerate these pre-cooled atoms through the DP tubes into the science cell, following a trajectory analogous to that of a projectile. For atoms to successfully traverse both stages of the DP tubes, they must possess a sufficiently high initial longitudinal velocity to complete the passage, yet low enough to remain within the capture range of the 3D MOT.

### v2

@@ I Introduction
Ultracold polar molecules have emerged as a powerful platform for studying ultracold chemistry, testing fundamental symmetry, and exploring strongly correlated many-body systems with strong dipole-dipole interactions [Carr2009, Ye2017]. Among bialkali polar molecules, 6Li87Rb ground-state molecules stand out as they exhibit the second-largest electric dipole moment [yin2024stirap]. This makes it a promising candidate for realizing and probing novel quantum phases, such as the long-sought dipolar Bardeen-Cooper-Schrieffer to Bose-Einstein condensate (BCS-BEC) crossover [deng2023]. So far, several labs have successfully produced degenerate mixtures of fermionic lithium and bosonic rubidium atoms [silber2005, taglieber2008] and observed Feshbach resonances between lithium and rubidium atoms [deh2008, deh2010]. Ultracold 6Li87Rb polar molecules remain elusive in experiments.
@@
Next-generation experiments with ultracold dual-species atoms and polar molecules require shorter cycle times and miniaturized system volume, thereby demanding a compact dual-species cold-atom source with high atomic flux. In the early days, achieving high flux often necessitated complex setups with multiple bulky vacuum chambers, especially for high-melting-point elements such as lithium and sodium [phillips1982Na, joffe1993Na, mewes1999_Li, hadzibabic2002, taglieber2006, marti2010LiRb, ridinger2011LiK, park2012, hopkins2016, wu2017LiK, ilzhofer2018, ma2019, Hulet2020_Li_QGM, warner2021, vayninger2025]. Those elements typically require long Zeeman slowers to decelerate fast atomic beams from hot ovens. While compact, two-dimensional magneto-optical traps (2D MOTs) have relatively low capture velocities, therefore tend to work more efficiently for low-melting-point elements such as rubidium and cesium [dieckmann1998a, lam2020, chenLithiumcesiumSlowBeam2021, pur2023, sutradhar2023, matthies2024]. In 2009, Tiecke et al. demonstrated a compact lithium 2D MOT using permanent magnets, achieving loading rates comparable to those of Zeeman slowers [tiecke2009]. Recently, a modularized lithium quantum gas setup based on a 2D MOT has been developed in Heidelberg and Munich [hammel2025, jain2025_unirand_li], which further enhances the loading rate by employing high-power and top-hat cooling beams [hammel2025]. Combining a 2D MOT with a short-distance Zeeman slower can substantially enhance its capture velocity and thereby loading rates. This has been demonstrated for sodium, strontium, and ytterbium atoms [lamporesi2013, li2020, li2023, nosske2017, wodey2021], but not for lithium atoms yet. Moreover, integrating high- and low-melting-point elements into a single compact system while maintaining high flux for both species remains an experimental challenge.
@@
In this work, we present a compact and high-flux dual-species cold-atom source that overcomes these challenges by combining a rubidium 2D MOT in series with a lithium 2D MOT, enhanced by a short-distance Zeeman slower for lithium. As a result, the 3D MOT loading rate of lithium atoms reaches a record value that is 44 times higher than that without the Zeeman slowing light. The flux of rubidium also reaches a reasonably high value. Meanwhile, the vacuum chamber is so small that it can be mounted on a movable breadboard [hammel2025]. Thereby, we can smoothly extract it from the magnetic coils and the forest of optics for laser beam alignment or maintenance, then push it back to the original position. Thanks to the in-series 2D MOTs configuration, the compact vacuum system also allows more than 270 degrees wide optical access to the science cell in the horizontal plane, providing an excellent starting point for future experiments.
@@ II.1 Vacuum system and general setup
Our vacuum system integrates Rb and Li two-dimensional magneto-optical traps (2D MOTs) arranged in series inside a high-vacuum (HV) region. These sources generate high-flux pre-cooled atomic beams that are further collimated by push beams and captured in a dual-species 3D MOT located in an ultra-high-vacuum (UHV) science cell, as shown in Fig. 1. Two differential pumping channels (DP) maintain a pressure drop of more than three orders of magnitude between the HV and UHV regions while preserving large solid angles for both beams (see Appendix A). Placing both 2D MOTs on the same side of the main chamber maximizes optical and microwave access to the science cell for manipulation of ultracold atoms and molecules [biswas2025_nak_microwave]. During daily operation, we measure a pressure on the order of $1\times 10^{-9}\ \mathrm{mbar}$ in the Li octagon chamber and $3\times 10^{-11}\ \mathrm{mbar}$ in the cubic science chamber. This pressure is consistent with the measured vacuum-limited lifetime exceeding 50s50\ $\mathrm${s} in the science chamber.
The Rb 2D MOT loads atoms directly from the background vapor at a pressure of about $10^{-8}\ \mathrm{mbar}$ and provides an atomic flux of roughly $10^{10}\ \mathrm{atoms/s}$ in a $2\mathrm{D}^{+}$ configuration. In contrast, $^{6}\mathrm{Li}$ has an extremely low vapor pressure at room temperature ($\sim 10^{-20}\ \mathrm{mbar}$[gehm2003] and therefore requires an oven-based source. We place the high-vapor-pressure, low-divergence Rb 2D MOT at the upper stream of the vacuum system, while the low-vapor-pressure, high-divergence Li 2D MOT is located in the middle, with the crucial UHV science chamber further downstream. The Li 2D MOT chamber also serves as an additional differential pumping stage between the Rb 2D MOT region and the science chamber. This layout facilitates a natural HV-to-UHV transition, allowing for the independent optimization of the spatially separated Rb and Li 2D MOTs, thereby enabling the simultaneous high-flux loading of both species into the 3D MOT.
@@ II.2 Source and Ovens
The Rb source (CAS: 7440-17-7, CAT: 044214.03, Thermo Fisher Scientific Inc.) consists of 1 gram of natural isotopic abundance rubidium ($\eta_{{}^{87}\mathrm{Rb}}=27.83(2)\%$ and $\eta_{{}^{85}\mathrm{Rb}}=72.17(2)\%$ [daniela.steck2023, daniela.steck2023a]), sealed in a glass ampoule to prevent oxidation. We designed the Rb oven using a copper tube with a specialized opening that serves as a gasket (LIMIT VACUUM, Beijing). This assembly is sealed with a custom flange and connected to the Rb chamber via an angle valve, allowing for a controlled and gentle introduction of the atomic vapor. To activate the Rb source, we mechanically crush the copper tube, breaking the enclosed ampoule during the final stage of the vacuum baking process.
@@ II.3.1 Lithium laser system
The $^{6}\text{Li}-D_{2}$ transition possesses a natural linewidth of $\Gamma_{\mathrm{Li}}/2\pi\approx 5.87\ \mathrm{MHz}$, while the hyperfine splitting between the upper manifold $|2^{2}\text{P}_{3/2};F^{\prime}=1/2,3/2,5/2\rangle$ is merely 4.4MHz4.4\ $\mathrm${MHz} [das2007, gehm2003]. While driving the cycling transition $|2^{2}\text{S}_{1/2};F=3/2\rangle\to|2^{2}\text{P}_{3/2};F^{\prime}=5/2\rangle$ for laser cooling, the transitions to the $|2^{2}\text{P}_{3/2};F^{\prime}=1/2,3/2\rangle$ states are simultaneously driven. Consequently, no convenient fully closed cooling transitions can be utilized here [tiecke2009].
@@
Figure 2: Laser frequency scheme (not to scale). Left: $^{6}\mathrm{Li}$ $D_{2}$ transition from the $2^{2}\mathrm{S}_{1/2}$ state to the $2^{2}\mathrm{P}_{3/2}$ manifold. Red arrows indicate the corresponding laser frequencies, and $\delta^{\mathrm{Li}}_{i}$ denote the detunings of the individual beams with respect to the upper manifold. Transition frequencies are taken from [gehm2003]. Right: $^{87}\mathrm{Rb}$ $D_{2}$ transition from the $5^{2}\mathrm{S}_{1/2}$ state to the $5^{2}\mathrm{P}_{3/2}$ state. Red arrows indicate the corresponding laser frequency, and $\delta^{\mathrm{Rb}}_{i}$ denote the detunings of the individual beams with respect to the corresponding transition. Transition frequencies are taken from [daniela.steck2023].
@@ II.3.2 Rubidium Laser System
The laser frequency scheme for laser cooling of $^{87}\text{Rb}$ is shown in Fig. 2, following many of the well-established Rb cooling schemes, for example [wei2020]. In contrast to $^{6}\text{Li}$, the $^{87}\text{Rb}\ D_{2}$ transition exhibits a natural linewidth of $\Gamma_{\mathrm{Rb}}/2\pi\approx 6.07\ \mathrm{MHz}$, with hyperfine splittings in the $5^{2}\mathrm{P}_{3/2}$ states ranging from tens to hundreds of $\mathrm{MHz}$ [daniela.steck2023, ye1996]. These well-resolved energy separations significantly simplify the implementation of laser cooling. For our cooling scheme, we utilize the closed $|5^{2}\mathrm{S}_{1/2};F=2\rangle\to|5^{2}\mathrm{P}_{3/2};F^{\prime}=3\rangle$ transition as the primary cooling cycle, supplemented by a repumping beam resonant with the $|5^{2}\mathrm{S}_{1/2};F=1\rangle\to|5^{2}\mathrm{P}_{3/2};F^{\prime}=2\rangle$ transition to recover atoms that decay out of the cooling cycle. Analogous to our lithium configuration, we implement a push beam, red-detuned from the $|5^{2}\mathrm{S}_{1/2};F=2\rangle\to|5^{2}\mathrm{P}_{3/2};F^{\prime}=3\rangle$ transition, to enhance the directed atomic flux along the axial direction.
@@ II.5 Li short-distance Zeeman slowing design
The simulations were conducted based on the simplified models outlined in [tiecke2009, lamporesi2013, nosske2017]. The total atomic flux emitted through the oven aperture (area $A$) is given by:
@@
This formula is a good approximation for the capture velocity $v_{c}\ll\alpha$[tiecke2009]. Here, $a_{6}$ is the isotopic abundance of $^{6}\text{Li}$, $\alpha=\sqrt{2k_{B}T/m}$ is the most probable velocity, and $\Omega_{c}$ is the solid angle subtended by the 2D MOT region. The oven temperature $T$ determines the velocity distribution and the saturated vapor pressure $p_{s}$, which is derived from the Clausius-Clapeyron relation. Further details can be found in [tiecke2009].
Figure 4: Simulation of the short-distance Li Zeeman slowing and 2D MOT. (a) Velocity distribution of $10^{6}$ Li atoms with and without Zeeman slowing. The Zeeman slower reshapes the initial distribution by decelerating atoms in the blue region into the orange peak, bringing them below the 2D MOT capture velocity $v_{c}$. Inset: velocity deformation caused by Zeeman slowing (figure inspired by [lamporesi2013]). (b) Numerical simulations of atomic trajectories along the applied magnetic field. The 2D MOT beam region is shown in orange. The light gray area indicates the capture velocity $v_{c}$ of the 2D MOT.
@@
To minimize contamination of the glass viewport caused by lithium’s high sticking coefficient and small mass, the Zeeman beam is redirected using a gold-coated in-vacuum mirror mounted 300mm300\ $\mathrm${mm} above the center of the octagon chamber on the top of a so-called double-T structure (compare Fig. 1). However, we still observe some contamination after several months of operation, although the ZS performance is barely affected. An alternative approach would be to heat a viewport made of sapphire crystal positioned directly against the atomic flux, as demonstrated by Nosske et al. [nosske2017]. Their implementation also showed degradation in optical transmission over a period of several months. The ZS beam is far-detuned (approximately $85\Gamma$) to take advantage of the positive gradient region of the magnetic field ($\partial B_{y}/\partial z>0$), minimizing interference with the 2D MOT operation. In contrast to conventional ZS designs [phillips1982Na], our system employs a quadrupole magnetic field oriented transverse to the atomic beam trajectory. This configuration demands circularly polarized light relative to the local magnetic field vector. Consequently, when using a linearly polarized ZS beam (as in our setup), only the $\sigma^{-}$ polarization component effectively contributes to the slowing process, reducing the practical intensity of the ZS beam by 50% [nosske2017]. Along one dimension, the radiation force acting on the atoms is described by:
@@
Here, $\delta_{0}$ is the zero-field detuning, $v$ is the atomic velocity, $B$ is the magnetic field strength, and the factor $0<\eta_{B}<1$ accounts for shielding effects of the metal chamber to the magnetic field[lamporesi2013].
@@ II.6 Push Beam
The atom cloud generated by the 2D MOT forms a needle-shaped distribution aligned along the longitudinal axis. This two-dimensional cooling process produces slow atomic beams with velocities typically below 30m/s30\ $\mathrm${m/s} [dieckmann1998a]. The push beam serves to accelerate these pre-cooled atoms through the DP tubes into the science cell, following a trajectory analogous to that of a projectile. For atoms to successfully traverse both stages of the DP tubes, they must possess a sufficiently high initial longitudinal velocity to complete the passage, yet low enough to remain within the capture range of the 3D MOT.